(Wallace)

Bell Industries sells samsung phones for CDN$100. The unit variable cost per phone is $50 plus a selling commission of 10%. Fixed manufacturing costs total $1,250 per month, while fixed selling and administrative costs total $2,500.

A. What is the contribution margin per phone?
Selling Price - $100.00
Variable Cost = CAD$50 + 10%($100)
Total Variable Cost = $50+10
Total Variable Cost = $60.00
Contribution Margin = Selling price - Variable Cost
= $100-60
= $40

B. What is the break even point ?

BEP = TFC /Unit Contribution
= 2500+1250 /40
= 94 ( Rounded off to nearest whole #)

C. How much phone must be sold for us to meet a target $7500?

Sales ( Unit) = TNI+TFC/Contribution Margin
= $3750+$7500/40
= $11250/40
= 281 phones

Are these answers correct/wrong, if so. Comment on answers
Answer #1

Answer A

Selling Price - $100.00(per unit)
Variable Cost = CAD$50 + 10%($100)(per unit)
Total Variable Cost = $50+10(per unit)
Total Variable Cost = $60.00(per unit)
Contribution Margin (per unit)= Selling price - Variable Cost
= $100-60
= $40

The answer is perfectly correct and the formula applied is very well written

Answer B

BEP = TFC /Unit Contribution
= 2500+1250 /40
= 94 (93.75 =94

Thus answer is completely correct

Answer C

Units at Target Profit = Fixed Costs + Target Income
Contribution Margin per unit

Sales ( Unit) = TNI+TFC/Contribution Margin
= $3750+$7500/40
= $11250/40
= 281 phones

The answer is correct and when we determine units at target point we assume it is necessary to earn particular income and assuming it to be a fixed cost (in a way )and then we determine BEP to calculate the number of units to be sold .

All the answers are correct and KUDOS to you !